Determine the margin of error for the confidence interval for the proportion: 0.509 <u<0.845 The margin of error is
Answer #1

Solution :

Point estimate = \hat p = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = \hat p = (0.509+0.845) / 2=0.677

Margin of error = E = Upper confidence interval - \hat p = 0.845-0.677=0.168

Margin of error = E =0.168